Ignition of MgNH4PO4·6H2O produces NH3, H2O and magnesium pyrophosphate, Mg2P2O7. Complete the balance equation for this reaction. If 5.00 g of MgNH4PO4·6H2O are ignited, how many grams of Mg2P2O7 would be formed?
1
Expert's answer
2020-09-16T06:50:16-0400

2MgNH4PO4·6H2O + O2 = 2NH3 + 7H2O + Mg2P2O7

M (MgNH4PO4·6H2O) = 245 g/mol

M (Mg2P2O7) = 126.6 g/mol

n (MgNH4PO4·6H2O) = m/M = 5/254=0.02 mol

n (Mg2P2O7) = 1/2 x n (MgNH4PO4·6H2O) = 0.02/2 = 0.01 mol

m (Mg2P2O7) = n x M = 0.01 x 126.6 = 1.3 g